# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ 5 Experiments
Results and Discussion for Q3. To empirically investigate the impact of width $d$ (i.e., the number of hidden units) on the generalization gap, we conduct additional experiments using a 5-layer GCN equipped with a normalized graph filter. The experiments specifically involve a comparison between a 5-layer GCN configured with a width of $2d$ for each layer and the previously studied model with $d$ width ($d=32$), as illustrated in Figure 3. This setup allows for a direct comparison under varying network configurations, providing insights into how changes in the number of hidden units influence the generalization gap. As demonstrated in Figure 3, across all the datasets examined, a $d$-width GCN consistently exhibits smaller generalization gaps compared to one with a $2d$-width. This observation is in harmony with our theoretical explanation presented after Theorem 1, that is, the factor $B$ (i.e., the upper bound of 2-norm of the parameters $\{\mathbf{W}^{(1)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$) directly influences factors $\kappa_{1}$ and $\kappa_{2}$ in the upper bound of the generalization gap.
@@ 6 Conclusion and Further Remarks
This paper explores the generalization of deep GCNs by providing an upper bound on their generalization gap. Our generalization bound is obtained based on the algorithmic stability of deep GCNs trained by the SGD algorithm. Our analysis demonstrates that the algorithmic stability of deep GCNs is contingent upon two factors: the largest absolute eigenvalue (or maximum singular value) of graph filter operators and the number of layers utilized. In particular, if the aforementioned eigenvalue (or singular value) remains invariant regardless of changes in the graph size, deep GCNs exhibit robust uniform stability, resulting in an enhanced generalization capability. Additionally, our results suggest that a greater number of layers can increase the generalization gap and subsequently degrade the performance of deep GCNs. This provides guidance for designing well-performing deep GCNs with a proper number of layers [56]. Most importantly, the result of single-layer GCNs in [36] can be regarded as a special case of our results in deep GCNs without hidden layers.
While our study is primarily focused on exploring the fundamental principles of generalizability and stability in the context of a simple deep GCN model framework, it can offer preliminary insights into several pressing issues that are the subject of recent attention in the GNN domain. These include: i) the over-smoothing problem, which stands as a pivotal challenge in the development of deep GNNs [57, 58], and ii) the design of advanced GNNs tailored for heterophilic graphs, characterized by nodes whose labels significantly diverge from those of their neighbors [59, 60]. Some further remarks on these two issues are as follows:
  • 1.
@@
    We note that, given a trivial deep GCN model characterized by over-smoothed node embeddings (which typically result in significant training errors), our theoretical upper bound still holds, that is, for a given graph filter, an increase of layers could potentially increase this upper bound in a probabilistic sense. This also motivates the exploration of advanced deep GCN models that incorporate mechanisms to counteract over-smoothing, like the skip connection trick used in GCNII [61] and its follow-up works. This observation encourages the investigation of more sophisticated deep GCN models that employ strategies to mitigate over-smoothing effects, such as the implementation of skip connections, a technique exemplified by GCNII and its subsequent developments. In both theory and practice, reducing the maximum absolute eigenvalue of graph filter operators is achievable through the strategic implementation of skip connections across layers, which can potentially reduce the generalization gap. From this perspective, we anticipate that our findings will inspire further studies into advanced deep GCN structures, especially those designed to mitigate the over-smoothing issue, offering a new direction for both theoretical exploration and practical application in advanced deep GCN architectures.
  • 2.
@@
    Expanding our theoretical insights to include specific models tailored for heterophily graphs is valuable but requires deliberate effort. This involves assessing the impact of the homophily/heterophily ratio on the input graph signal, and incorporating this ratio into the upper bound estimation. It is important to clarify that, although our current empirical study considers two types of low-pass filters, the scope of our theoretical findings is not restricted to low-pass scenarios alone. To ensure a consistent and fair empirical evaluation, as demonstrated in [36], we utilized the benchmark graph datasets (Cora, Citeseer, Pubmed) known for their homophilic properties in node classification tasks. However, for analyses involving high-pass filters, it would be appropriate to engage with benchmark datasets representing heterophily graphs (such as Texas, Wisconsin, Cornell, etc.). We refer the readers interested in delving deeper into this topic to the recent work [46], in which the authors use analytical tools from statistical physics and random matrix theory to precisely characterize generalization in simple graph convolution networks on the contextual stochastic block model (CSBM). This research, though based on specific assumptions on the graph signal, can inspire further refinements in our theoretical framework, outcomes, and methodologies, taking into account unique graph signal characteristics (e.g., homophily/heterophily) and model complexities (e.g., low-pass/high-pass filters, depth and width of network architecture).
@@
In terms of future research directions, it would be valuable to extend the theoretical analysis presented in this study to encompass other commonly used learning algorithms in graph neural networks, moving beyond the scope of SGD. Moreover, our theoretical results offer insights that can inform the exploration of various strategies to enhance the generalization capability of deep graph neural networks. This could involve investigating the efficacy of regularization techniques, conducting advanced network architecture searches, or developing adaptive graph filters. Additionally, a significant area for future investigation is to establish the potential connection between the model’s stability and generalization, and the issues of over-smoothing and over-squashing encountered in deep graph neural networks. Understanding these interrelationships can potentially contribute to the development of novel techniques and algorithms that address these challenges and improve the overall effectiveness of deep graph neural networks in dealing with more complex tasks.
Appendix A Proofs

### v2

@@ 5 Experiments
Results and Discussion for Q3. To empirically investigate the impact of width $d$ (i.e., the number of hidden units) on the generalization gap, we conduct additional experiments using a 5-layer GCN equipped with a normalized graph filter. The experiments specifically involve a comparison between a 5-layer GCN configured with a width of $2d$ for each layer and the previously studied model with $d$ width ($d=32$), as illustrated in Fig. 3. This setup allows for a direct comparison under varying network configurations, providing insights into how changes in the number of hidden units influence the generalization gap. As demonstrated in Fig. 3, across all the datasets examined, a $d$-width GCN consistently exhibits smaller generalization gaps compared to one with a $2d$-width. This observation is in harmony with our theoretical explanation presented after Theorem 1, that is, the factor $B$ (i.e., the upper bound of 2-norm of the parameters $\{\mathbf{W}^{(1)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$) directly influences factors $\kappa_{1}$ and $\kappa_{2}$ in the upper bound of the generalization gap.
6 Theoretical Implications
Our work establishes a theoretical framework for analyzing the generalization gap of traditional deep GCNs, which further provides insights into extending the analysis to other classes of graph neural networks, including Graph Transformers. As illustrative examples, we briefly discuss how the theoretical proof methodology developed in our framework can be applied to GCNII and Graph Transformer, which are representative models of more advanced GNNs, thereby demonstrating the broader applicability of our theoretical framework.
6.1 Extension to GCNII
With input features $\mathbf{X}^{(0)}=\mathbf{X}\in\mathbb{R}^{N\times d}$, GCNII defines its $k$-th layer as
$\mathbf{X}^{(k)}=\sigma\bigg($ $\Big((1-a_{k})g(\mathbf{L})\mathbf{X}^{(k-1)}+a_{k}\mathbf{X}^{(0)}\Big)\cdot\Big((1-b_{k})\mathbf{I}_{d}+b_{k}\mathbf{W}^{(k)}\Big)\bigg),$
for $k=1,2,\ldots,K$, where $a_{k},b_{k}\in(0,1)$ are two hyperparameters, $\mathbf{X}^{(k)}$ is the output feature matrix of the $k$-th layer, $\mathbf{W}^{(k)}$ is the trained parameter matrix specific to the $k$-th layer, graph filter $g(\mathbf{L})=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-1/2}$, and $\mathbf{I}_{d}$ is the $d\times d$ identity matrix. The output for node $\mathbf{x}$ is
$f(\mathbf{x}|\theta)=\sigma\left(\boldsymbol{\delta}_{\mathbf{x}}^{\top}\left((1-a_{K+1})g(\mathbf{L})\mathbf{X}^{(K)}+a_{K+1}\mathbf{X}^{(0)}\right)\mathbf{w}\right),$
where $\theta=\{\mathbf{W}^{(1)},\mathbf{W}^{(2)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$ (all trainable parameters, with $\mathbf{w}\in\mathbb{R}^{d}$ the output layer parameter); $\boldsymbol{\delta}_{\mathbf{x}}\in\mathbb{R}^{N}$ is the indicator vector for node $\mathbf{x}$; $a_{K+1}\in(0,1)$ is a hyperparameter for the output layer residual connection. Let $\theta_{t}$ and $\theta^{\prime}_{t}$ be the learned parameters of two GCNs trained on $\mathcal{S}$ and $\mathcal{S}^{i}$ using SGD in the $t$-th iteration with $\theta_{0}=\theta_{0}^{\prime}$, and $\triangle\theta_{t}:=\theta_{t}-\theta^{\prime}_{t}$.
For each layer $k$, the perturbation of layer outputs $\|\triangle\mathbf{X}^{(k)}\|_{F}=\|\mathbf{X}^{(k)}-\mathbf{X}^{(k)^{\prime}}\|_{F}$ satisfies the recursive bound:
$\|\triangle\mathbf{X}^{(k)}\|_{F}\leq c_{1}^{(k)}\|\triangle\mathbf{X}^{(k-1)}\|_{F}+c_{2}^{(k)}\|\triangle\mathbf{W}^{(k)}\|_{2},$ (19)
where $c_{1}^{(k)}=(1-a_{k})(1-b_{k}+b_{k}B)\alpha_{\sigma}C_{g}$ and $c_{2}^{(k)}=\alpha_{\sigma}b_{k}\big((1-a_{k})C_{g}B_{\mathbf{X}}^{(k-1)}+a_{k}C_{\mathbf{X}}\big)$ with $B_{\mathbf{X}}^{(k-1)}$ the bound of $\|\mathbf{X}^{(k-1)}\|_{F}$ (see (A.22) in the Appendix E). The first term on the right side of the iterative formula captures propagation of perturbations from the previous layer, while the second term captures perturbation from $\mathbf{W}^{(k)}$.
By induction, it yields that
$\|\triangle\mathbf{X}^{(k)}\|_{F}\leq e^{(k)}(\sum_{j=1}^{k}\|\triangle\mathbf{W}^{(k)}\|_{2}),$ (20)
where $e^{(k)}=\max\{c_{1}^{(k)}e^{(k-1)},c_{2}^{(k)}\}$ with $e^{(0)}=0$. We provide the proof of (19) and (20) in Appendix E. Then, combining layer-wise bounds and using the Lipschitz property of $\sigma$, one can have the output perturbation $|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|$ bounded by the total parameter perturbation $\|\Delta\theta\|_{*}=\sum\limits_{j=1}^{K}\|\mathbf{W}^{(j)}-\mathbf{W}^{(j)^{\prime}}\|_{2}+\|\mathbf{w}-\mathbf{w}^{\prime}\|_{2}$ (see Appendix E for technical details) as
$|f(\mathbf{x}|\theta)-f(\mathbf{x}|\theta^{\prime})|\leq\alpha_{\sigma}\cdot\varrho\|\triangle\theta\|_{*},$ (21)
where $\varrho=\max\Big\{(1-a_{K+1})BC_{g}\cdot e^{(K)},(1-a_{K+1})C_{g}B_{\mathbf{X}}^{(K)}+a_{K+1}C_{\mathbf{X}}\Big\}$. Then,
$\Big|\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y},y)\big]-\mathbb{E}_{\mathcal{A}}\big[\ell(\hat{y}^{\prime},y)\big]\Big|=\Big|\mathbb{E}_{\mathcal{A}}\big[\ell\big(f(\mathbf{x}|\theta_{T}),y\big)-\ell\big(f(\mathbf{x}|\theta^{\prime}_{T}),y\big)\big]\Big|\leq\alpha_{\ell}\mathbb{E}_{\mathcal{A}}\Big[\big|f(\mathbf{x}|\theta_{T})-f(\mathbf{x}|\theta^{\prime}_{T})\big|\Big]\leq\varrho\alpha_{\ell}\cdot\mathbb{E}_{\mathcal{A}}\big[\|\triangle\theta_{T}\|_{*}\big].$
This implies that the stability of $\mathcal{A}_{\mathcal{S}}$ for GCNII has a bound
$\mu_{m}\leq\frac{\varrho\alpha_{\ell}}{2}\sup_{\mathcal{S}}\Big\{\mathbb{E}_{\mathcal{A}}[\|\triangle\theta_{T}\|_{*}]\Big\}.$
Note that when $a_{k}=0,b_{k}=1$ for all $k$, GCNII degenerates into the traditional GCN, we have $\varrho=B^{K}\alpha_{\sigma}^{K}C_{g}^{K+1}C_{\mathbf{X}}$, and thus
$\mu_{m}\leq\frac{\alpha_{\ell}B^{K}\alpha_{\sigma}^{K}C_{g}^{K+1}C_{\mathbf{X}}}{2}\sup_{\mathcal{S}}\Big\{\mathbb{E}_{\mathcal{A}}[\|\triangle\theta_{T}\|_{*}]\Big\},$
which is consistent with (14).
To further bound $\|\triangle\theta_{T}\|_{*}$, the crucial step is to bound the perturbation of the gradient of $f(\mathbf{x}|\theta)$ with respect to the parameters $\theta=\{\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{K},\mathbf{w}\}$ and obtain the result similar to Lemma 7 in Appendix C, which can be achieved by following the technique in our paper. Here, we provide the result for $\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\|_{F}$:
$\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\|_{F}\leq\Big(\nu_{\sigma}\varrho\cdot\big((1-a_{K+1})C_{g}B_{\mathbf{X}}^{(K)}+a_{K+1}C_{\mathbf{X}}\big)+\alpha_{\sigma}\cdot(1-a_{K+1})C_{g}e^{(K)}\Big)\cdot\|\triangle\theta\|_{*},$ (22)
where $\varrho=\max\Big\{(1-a_{K+1})BC_{g}\cdot e^{(K)},(1-a_{K+1})C_{g}B_{\mathbf{X}}^{(K)}+a_{K+1}C_{\mathbf{X}}\Big\}$. Note that when $a_{k}=0,b_{k}=1$ for all $k$, GCNII degenerates into the traditional GCN, we have $\varrho=B^{K}\alpha_{\sigma}^{K}C_{g}^{K+1}C_{\mathbf{X}}$, $B_{\mathbf{X}}^{(K)}=B^{K}\alpha_{\sigma}^{K}C_{g}^{K}C_{\mathbf{X}}$ and $e^{(K)}=B^{K-1}\alpha_{\sigma}^{K}C_{g}^{K}C_{\mathbf{X}}$. At this point,
$\big\|\nabla_{\mathbf{w}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{w}}f(\mathbf{x}|\theta^{\prime})\big\|_{F}\leq\Big(\upsilon_{\sigma}B^{2K}\alpha^{2K}_{\sigma}C^{2K+2}_{g}C^{2}_{\mathbf{X}}+B^{K-1}\alpha^{K+1}_{\sigma}C^{K+1}_{g}C_{\mathbf{X}}\Big)\|\triangle\theta\|_{*},$
which is consistent with (A.10) in Appendix C. For the bound of $\|\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta)-\nabla_{\mathbf{W}^{(k)}}f(\mathbf{x}|\theta^{\prime})\|_{F}$, we refer the readers to the proof process of (A.27) in Appendix E.
Finally, these structured analysis results can lead to the results corresponding Lemma 3 and Lemma 4, and thus enable bounding the stability of GCNII.
6.2 Extension to Graph Transformer
To extend our theoretical framework to more complex models like Graph Transformer, the key is to bound the generalization gap of Graph Transformer by quantifying how perturbations in the training set (e.g., removing or replacing a node) propagate to changes in model outputs. Graph Transformer introduce new learnable parameters: query $(\mathbf{W}_{Q})$, key $(\mathbf{W}_{K})$, and value $(\mathbf{W}_{V})$ projection matrices, alongside attention scalers and feed-forward layers, for which a self-attention layer is defined GT as
$F\left(\mathbf{x}_{n}\right)=\mathbf{a}^{\top}\operatorname{Relu}\Big(\mathbf{W}_{O}\sum_{i\in\mathcal{T}^{n}}\mathbf{W}_{V}\mathbf{x}_{i}\cdot\operatorname{softmax}_{n}\big((\mathbf{W}_{K}\mathbf{x}_{i})^{\top}\mathbf{W}_{Q}\mathbf{x}_{n}\big)\Big),$
where $\mathbf{x}_{i}$ denotes features of node $i$, $\mathcal{T}^{n}$ is the set of nodes for the aggregation computing of node $n$, and $\operatorname{softmax}_{n}(h(i,n))=\exp(h(i,n))/\sum_{j\in\mathcal{T}^{n}}\exp(h(j,n)).$ Despite their architectural complexity (e.g., self-attention mechanisms, query/key/value projections), gradient decomposition still remains to be conducted via the product rule and chain rule, accounting for the propagation of attention-weight variations to the final output. Besides, a Lipschitz-type inequality for softmax may be critically needed, for which we claim that for $\mathbf{z}=(z_{1},z_{2},\dots,z_{p})$, $\mathbf{z}^{\prime}=(z^{\prime}_{1},z^{\prime}_{2},\dots,z^{\prime}_{p})$ with $\|\mathbf{z}-\mathbf{z}^{\prime}\|_{\infty}\leq 1$,
$\|\operatorname{softmax}(\mathbf{z})-\operatorname{softmax}(\mathbf{z}^{\prime})\|_{1}\leq 2e\|\mathbf{z}-\mathbf{z}^{\prime}\|_{\infty}.$ (23)
Actually, the proof is not hard to set up by straight forward boundedness and the mean value theorem of exponential functions (see the technical details in Appendix F).
For trainable parameters $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$, set the attention output is:
$F(\mathbf{x}_{n})=\mathbf{a}^{\top}\text{ReLu}\Big(\mathbf{W}_{O}\sum_{i\in\mathcal{T}^{n}}\mathbf{W}_{V}\mathbf{x}_{i}\cdot\operatorname{Attn}(\mathbf{x}_{n})_{i}\Big),$
where $S_{i,n}=(\mathbf{W}_{K}\mathbf{x}_{i})^{T}(\mathbf{W}_{Q}\mathbf{x}_{n})$ is the scaled dot-product score, $A_{i,n}=\operatorname{softmax}_{n}(S_{i,n})$ are attention weights, and $\operatorname{Attn}(\mathbf{x}_{n})=\sum_{i\in\mathcal{T}^{n}}\mathbf{W}_{V}\mathbf{x}_{i}\cdot A_{i,n}$ the attention output. Then the gradient decomposition with respect to $\mathbf{W}_{K}$ is given by
[…]
where $\mathbf{Z}=\mathbf{W}_{O}\cdot\text{Attn}(\mathbf{x}_{n})$, $\mathbf{A}=\{A_{i,n}\}$, and $\mathbf{S}=\{S_{i,n}\}$. Then calculating each item gives that
$\nabla_{\mathbf{W}_{K}}F(\mathbf{x}_{n})=$ $\mathbf{a}^{\top}\mathbb{I}_{\geq 0}(\mathbf{W}_{O}\cdot\text{Attn}(\mathbf{x}_{n}))\cdot\mathbf{W}_{Q}\cdot\mathbf{W}_{V}\cdot\left(\sum_{i\in\mathcal{T}^{n}}\mathbf{A}_{i,n}(\mathbf{x}_{i}-\bar{\mathbf{x}}_{n})\mathbf{x}_{i}^{\top}\right)\cdot(\mathbf{W}_{Q}\mathbf{x}_{n})^{\top}.$
By leveraging the Lipschitz continuity of the gradient with respect to its trainable parameters, it can lead to bounding the gradient perturbation in terms of the total parameter perturbation $\|\triangle\theta\|_{*}=\|\mathbf{W}_{K}-\mathbf{W}^{\prime}_{K}\|_{2}+\|\mathbf{W}_{V}-\mathbf{W}^{\prime}_{V}\|_{2}+\|\mathbf{W}_{O}-\mathbf{W}^{\prime}_{O}\|_{2}+\|\mathbf{W}_{Q}-\mathbf{W}^{\prime}_{Q}\|_{2}+\|\mathbf{a}-\mathbf{a}^{\prime}\|_{2}$ by
$\|\nabla_{\mathbf{W}_{K}}F(\mathbf{x}_{n}|\theta)-\nabla_{\mathbf{W}_{K}}F(\mathbf{x}_{n}|\theta^{\prime})\|_{2}\leq 2eK_{\max}B^{3}C_{\mathbf{X}}^{3}\|\Delta\theta\|_{*},$ (24)
where $K_{\max}\geq|\mathcal{T}^{n}|$ is the maximum neighborhood size, $B$ is the upper bound of weight matrices (technical details in Appendix F). It mirrors the Lemma 7 in our approach for deep GCNs, where we recursively decomposed gradients across layers (see Lemma 7). For Graph Transformer, similar recursive relations can be derived for attention layers, with additional terms capturing interactions between $\mathbf{W}_{Q}\mathbf{X},\mathbf{W}_{K}\mathbf{X},\mathbf{W}_{V}\mathbf{X}$. For GCNs, we bounded gradient variations using norms of graph filters and layer parameters (e.g., $\|g(\mathbf{L})\|_{2}$, $\|\mathbf{W}^{(k)}\|_{2}$). For Graph Transformer, this will be extended to: singular values of $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ (analogous to $C_{g}$ in GCNs), as they control the ”strength” of feature projections and Lipschitz constants of softmax and feed-forward activations (replacing $\alpha_{\sigma}$ for GCN activations, and leads to an analogous to Theorem 2 for deep GCNs.
@@ 7 Conclusion and Further Remarks
This paper explores the generalization of deep GCNs by providing an upper bound on their generalization gap. Our generalization bound is obtained based on the algorithmic stability of deep GCNs trained by the SGD algorithm. Our analysis demonstrates that the algorithmic stability of deep GCNs is contingent upon two factors: the largest absolute eigenvalue (or maximum singular value) of graph filter operators and the number of layers utilized. In particular, if the aforementioned eigenvalue (or singular value) remains invariant regardless of changes in the graph size, deep GCNs exhibit robust uniform stability, resulting in an enhanced generalization capability. Additionally, our results suggest that a greater number of layers can increase the generalization gap and subsequently degrade the performance of deep GCNs. This provides guidance for designing well-performing deep GCNs with a proper number of layers li2023deepergcn . Most importantly, the result of single-layer GCNs in 2019Stable_GCN can be regarded as a special case of our results in deep GCNs without hidden layers.
While our study is primarily focused on exploring the fundamental principles of generalizability and stability in the context of a simple deep GCN model framework, the theoretical insights obtained here can also offer preliminary perspectives on several research topics that have drawn increasing attention in the graph neural network community. These include, among others, the over-smoothing problem in deep architectures rusch2023survey , chen2022bag , the design of models tailored for heterophilic graphs zheng2022graph , zhu2023heterophily , and the emerging topic of graph out-of-distribution (OOD) generalization li2022ood , li2025out . Our theoretical study can provide potential hints toward these directions, but more fine-grained and comprehensive work is still needed to fully address them. Below, we elaborate on these aspects in turn, aiming to clarify their conceptual connections with our work, outline possible directions for extending our theoretical framework, and highlight three open and challenging questions that can serve as seeds for future exploration.
@@
How can the impact of over-smoothing in deep GCNs be mitigated? We first note that, given a trivial deep GCN model characterized by over-smoothed node embeddings (which typically result in significant training errors), our theoretical upper bound still holds — that is, for a given graph filter, an increase in layers could potentially increase this upper bound in a probabilistic sense. This also motivates the exploration of advanced deep GCN models that incorporate mechanisms to counteract over-smoothing, such as the skip connection technique used in GCNII chen2020simple and its follow-up works. As detailed in Section 5, our theoretical results can in fact be extended to the setting of GCNII, thereby providing analytical support for architectures that integrate skip connections. In both theory and practice, reducing the maximum absolute eigenvalue of graph filter operators is achievable through the strategic implementation of skip connections across layers, which can potentially reduce the generalization gap. From this perspective, our findings may inspire further studies into sophisticated deep GCN architectures designed to mitigate over-smoothing, offering a promising direction for both theoretical and practical advancements.
@@
What is the role of heterophily in GCN generalization? It is also valuable to consider extending our theoretical analysis to models specifically designed for heterophilic graphs, where nodes often connect to neighbors with dissimilar labels. This would require incorporating the homophily/heterophily ratio of the input graph signal into the upper bound estimation, thereby capturing how graph signal characteristics influence generalization. Although our empirical study here considers two types of low-pass filters on homophilic benchmark datasets (Cora, Citeseer, Pubmed), our theoretical framework is not restricted to low-pass scenarios alone. As remarked in Section 4.2, the analysis framework is in principle applicable to a broader range of filtering schemes; however, the derivations in our proofs do not explicitly examine the impact of specific quantities such as the homophily/heterophily ratio, leaving this as an open aspect for further refinement. To ensure a consistent and fair empirical evaluation, as demonstrated in 2019Stable_GCN , we adopt homophilic datasets that are standard in prior stability and generalization analyses of GCNs. For analyses involving high-pass filters, it would be appropriate to engage with heterophilic benchmark datasets (e.g., Texas, Wisconsin, Cornell). Relevant to this discussion is the recent work shi2024homophily , which employs analytical tools from statistical physics and random matrix theory to precisely characterize generalization in simple GCNs on the contextual stochastic block model (CSBM). Such studies, although based on specific graph signal assumptions, could inspire refinements to our theoretical framework by jointly considering graph signal characteristics (homophily/heterophily) and model complexities (filter types, depth, and width).
Can insights from in-distribution generalization inform OOD generalization? Beyond the above considerations, another relevant line of research that has recently attracted considerable attention is graph out-of-distribution (OOD) generalization li2022ood , li2025out . It is worth clarifying that the problem setting and theoretical assumptions in OOD generalization are distinct from those in the in-distribution generalization framework considered in this work. In-distribution generalization focuses on scenarios where both training and test data are drawn from the same underlying distribution, enabling rigorous analysis under well-defined stochastic assumptions, such as those adopted in our stability-based framework. In contrast, OOD generalization addresses cases involving distribution shifts, which often require additional modeling principles (e.g., invariance to spurious correlations, causal structure modeling, or domain adaptation techniques) and seek performance guarantees that hold across domains. Despite these differences, the two areas can be mutually beneficial: in-distribution analyses, such as our characterization of bias–variance trade-offs and the influence of spectral properties of graph filters on generalization, may offer insights for developing more OOD-robust architectures; conversely, OOD-oriented approaches, such as invariant risk minimization or causal subgraph intervention, may inspire new regularization schemes or architectural components that also enhance in-distribution performance. Related to this discussion, the authors in baranwal2021graph analyze a one-layer GCN trained on the CSBM via logistic regression, providing theoretical insights into improved linear separability and out-of-distribution generalization in semi-supervised node classification. Extending the current stability-based framework to accommodate mild forms of distribution shift thus presents an appealing research direction that could bridge these two lines of work and advance the understanding of generalization in graph neural networks.
Taken together, these discussions highlight that our theoretical framework, while developed under a specific in-distribution setting, has the potential to be extended and adapted to address a broader range of challenges in graph learning.
@@
Building on the above open questions, which outline core challenges for future exploration, it is also important to consider more concrete research directions and methodological extensions. For example, the theoretical analysis presented in this study could be extended to encompass other commonly used learning algorithms in graph neural networks, moving beyond the scope of SGD. Our theoretical results may also inform the exploration of strategies to enhance the generalization capability of deep graph neural networks, such as investigating the efficacy of regularization techniques, conducting advanced network architecture searches, or developing adaptive graph filters. In addition, establishing the potential connection between model stability, generalization, and the issues of over-smoothing and over-squashing represents another promising avenue. Understanding these interrelationships could contribute to the development of novel techniques and algorithms that address these challenges, thereby complementing the broader problem-oriented directions discussed above and improving the overall effectiveness of deep graph neural networks in dealing with more complex tasks.
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China (No. U21A20473, No. 62536006, No. 62172370). M. Li also acknowledged the support from the “Pioneer” and “Leading Goose” R&D Program of Zhejiang (No. 2024C03262). G. Yang acknowledged the support from the Opening Project of Guangdong Province Key Laboratory of Computational Science at the Sun Yat-sen University (No. 2024008). H. Feng was supported in part by the Research Grants Council of Hong Kong (Project no. CityU 11303821 ,and CityU 11315522). X. Zhuang was supported in part by the Research Grants Council of Hong Kong (Project no. CityU 11309122, CityU 11302023, CityU 11301224, and CityU 11300825). The authors also wish to thank Dr. Yi Wang (City University of Hong Kong, Hong Kong SAR, China) and Dr. Xianchen Zhou (National University of Defense Technology, China) for their insightful discussions and dedicated assistance with the experimental studies.
Appendix: Preliminaries